英語 (English)

基盤英語 (Basic English) (医・歯 ((医 a ・ (歯・口)A-1)1 年))

座喜純 · 准教授/ 大学院ソシオ・アーツ・アンド・サイエンス研究部 1 単位前期水 7 ・ 8

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

【授業の目的】The primary aim of this course is to improve the student’s basic ability to communicate in English in a variety of formal and informal situations. This course has many speaking and listening exercises that should increase the student’s ability to use English meaningfully on at least a basic level.

【授業の概要】Classes will be running in ENGLISH language only for more practices chance. There is no specific text book but there will be some other media depending on the ’Topics’ which will be selected up to date. Speaking out practices will take place during the classes.

【キーワード】 Basic English 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 The aim of this course is to improve the student’s ability to use basic

English to communicate in a variety of formal and informal situations. 【授業の計画】

1. Class Introduction

2. Self-Introduction

3. Clip 1

4. Clip 2

5. Clip 3

6. Clip 4

7. Clip 5

8. Clip 6

9. Clip 7

10. Clip 8

11. Clip 9

12. Clip 10

13. Clip 11

14. Clip 12

15. Review 【教科書】 [ ]There will be teacher handouts and text materials created by students.

【参考書等】 [参考資料]

【成績評価の方法】 Class presentations and written summaries of presentations, oral/written examinations and participation will be used for evaluation. Tentative evaluation scheme: Presentations=30% (3 X 10%), Examinations=20% (2 X 10%), Participation=50%.

【再試験の有無】 There is no make-up examination for this course.

【受講者へのメッセージ】 Class structure depends on the number of students enrolled. It may be possible to divide the class into smaller groups to facilitate conversation activities.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220825 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ 座喜 . 【備考】 This schedule is tentative and may be modified due to class size or other

factors.

English

Basic English (医・歯 ((医 a ・ (歯・口)A-1)1 年))

Jun Zaki · Associate Professor / Institute of Socio-Arts and Sciences 1 unit 前期水 7 ・ 8

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

Target⟩ The primary aim of this course is to improve the student’s basic ability to communicate in English in a variety of formal and informal situations. This course has many speaking and listening exercises that should increase the student’s ability to use English meaningfully on at least a basic level.

Outline⟩ Classes will be running in ENGLISH language only for more practices chance. There is no specific text book but there will be some other media depending on the ’Topics’ which will be selected up to date. Speaking out practices will take place during the classes.

�������⟩ Basic English Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ The aim of this course is to improve the student’s ability to use basic

English to communicate in a variety of formal and informal situations.

Schedule⟩

1. Class Introduction

2. Self-Introduction

3. Clip 1

4. Clip 2

5. Clip 3

6. Clip 4

7. Clip 5

8. Clip 6

9. Clip 7

10. Clip 8

11. Clip 9

12. Clip 10

13. Clip 11

14. Clip 12

15. Review Textbook⟩ [ ]There will be teacher handouts and text materials created by students.

Reference⟩ [参考資料]

Evaluation Criteria⟩ Class presentations and written summaries of presentations, oral/written examinations and participation will be used for evaluation. Tentative evaluation scheme: Presentations=30% (3 X 10%), Examinations=20% (2 X 10%), Participation=50%.

Re-evaluation⟩ There is no make-up examination for this course.

Message⟩ Class structure depends on the number of students enrolled. It may be possible to divide the class into smaller groups to facilitate conversation activities.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220825 Contact (Office-Hour, Room, E-mail)⟩

⇒ Zaki .

Note⟩ This schedule is tentative and may be modified due to class size or other factors.